# System-Embedded Diffusion Bridge Models

### v1

@@ 3 Background
where $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}$ is the score function and $\overline{\mathbf{w}}_{t}$ is the Wiener process with reversed time. The forward-reverse relationship ensures that the marginal distributions $p(\mathbf{x}_{t})$, defined by eq. 1 and eq. 2, remain consistent.
In standard SGM for image synthesis, $p(\mathbf{x}_{1})$ is typically chosen as an isotropic Gaussian. To generate samples from $p(\mathbf{x}_{0})$, one samples from $p(\mathbf{x}_{1})$ and solves eq. 2 using a time discretization scheme. This requires the unknown score function $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}$, which is generally approximated by a neural network $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ trained with the score-matching loss (Song et al., 2021b):
@@
The linear formulation in eq. 1 and eq. 2 is a special case of the broader SGM framework introduced by Song et al. (2021b), where drift and diffusion coefficients can be arbitrarily complex functions of $\mathbf{x}_{t}$ and $t$. This linear formulation, however, generalizes the well-established scalar case. Specifically, the Variance-Preserving (VP) SDE, which maintains constant variance for $\mathbf{x}_{t}$ across time, can be recovered with $\mathbf{F_{t}}=-\frac{1}{2}\beta_{t}\mathbf{I}$ and $\mathbf{G}_{t}=\sqrt{\beta_{t}}\mathbf{I}$. The Variance-Exploding (VE) SDE, which induces unbounded growth in the variance of $p(\mathbf{x}_{1})$ as $t\to 1$, is obtained with $\mathbf{F_{t}}=\mathbf{0}$ and $\mathbf{G}_{t}=\sqrt{\frac{\text{d}\sigma^{2}_{t}}{\text{d}t}}\mathbf{I}$. While these scalar matrices offer simplicity, they also raise the question of whether more sophisticated designs could yield additional benefits.
@@
where $\mathbf{y}\in\mathbb{R}^{m}$ is the measurement, $\boldsymbol{A}\in\mathbb{R}^{m\times d}$ is the system response matrix, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0}_{m},\mathbf{I}_{m\times m})$ is the measurement noise with covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{m\times m}$ and $p(\mathbf{y}\mid\mathbf{x})=\mathcal{\boldsymbol{N}}(\mathbf{A}\mathbf{x},\boldsymbol{\Sigma})$ is the forward model. We refer to finding the unknown $\mathbf{x}$ based on the provided $\mathbf{y}$ as solving a linear inverse problem with additive Gaussian noise (Tarantola, 2005). This formulation is central to many practical applications, such as signal processing, medical imaging, and remote sensing (Mueller and Siltanen, 2012), particularly in the case of noninvertible linear systems, where either $m\neq d$ or $\mathbf{A}$ is rank-deficient.
@@
Consider the simplest noiseless ($\boldsymbol{\Sigma}=\mathbf{0}_{m\times m}$) setting of eq. 4 with invertible $\mathbf{A}$. In this case, the original signal $\mathbf{x}$ can be fully recovered by simply using the inverse of $\mathbf{A}$, i.e., $\mathbf{x}=\mathbf{A}^{-1}\mathbf{y}$. However, this is only possible when $m=d$ and $\mathbf{A}$ is full-rank, which greatly limits the scope of potential inverse problems. In practice, the matrix $\mathbf{A}$ is typically not invertible, and noise arises naturally due to, e.g., imperfections in the measurement process.
@@
Many unsupervised approaches, assuming access to a specific noisy linear system (eq. 4), offer tailored solutions that improve performance under additional assumptions (Song et al., 2023a; Wang et al., 2023; Chung et al., 2023a; Garber and Tirer, 2024). In contrast, state-of-the-art (SOTA) diffusion bridges (Liu et al., 2023; Luo et al., 2023a; Yue et al., 2024; Zhou et al., 2024) make no assumptions about the underlying system during training, focusing on general mappings between arbitrary distributions. This can be suboptimal when the system is (even approximately) known. For instance, in image inpainting, these methods do not distinguish between the known range (unmasked) and the missing null part (masked), resulting in redundant noise in the range space, which is expected to be noiseless. In more complex scenarios, such redundancy could reduce efficiency, while ignoring system-specific information may hinder generalization. Thus, developing a supervised diffusion bridge tailored to a specific system remains an important yet unresolved challenge.
@@ 4 Method
Following eq. 4, it is clear that $\mathbf{y}\mid\mathbf{x}\sim\boldsymbol{\mathcal{N}}(\mathbf{A}\mathbf{x},\boldsymbol{\Sigma})$, i.e., for a given signal $\mathbf{x}$, a noisy linear system applies a linear transformation $\mathbf{A}$ and adds (possibly correlated) Gaussian noise to sample a measurement. On the other hand, given a clean sample $\mathbf{x}_{0}$, eq. 1 transforms it to $\mathbf{x}_{t}$ by applying a cascade of such noisy linear systems, which is a noisy linear system itself. That is, $\mathbf{x}_{t}\mid\mathbf{x}_{0}\sim\boldsymbol{\mathcal{N}}(\mathbf{H}_{t}\mathbf{x}_{0},\boldsymbol{\Sigma}_{t})$ for some time-dependent matrices $\mathbf{H}_{t},\boldsymbol{\Sigma}_{t}$. Hence, one may equivalently arrive at $\mathbf{y}\mid\mathbf{x}$ (or some linear transformation of $\mathbf{y}$) through an application of a series of noisy linear systems to $\mathbf{x}$ instead of just a single one. In the limit, this would imply the existence of a general linear SDE that performs such mapping.
@@
(Tivnan et al., 2025) Assume that $\mathbf{x}_{t}\mid\mathbf{x}_{0}$ evolves according to the linear forward process from eq. 1. Then, $\mathbf{x}_{t}\mid\mathbf{x}_{0}\sim\mathbf{\mathcal{N}}(\mathbf{H}_{t}\mathbf{x}_{0},\boldsymbol{\Sigma}_{t})$, where
@@
where $\boldsymbol{\Omega}_{t}$ is the Magnus expansion and the approximation becomes equality if, for all $s,s^{\prime}\in[0,t]$, $[\mathbf{F}_{s},\mathbf{F}_{s^{\prime}}]=0$, i.e., $\mathbf{F}_{s}$ and $\mathbf{F}_{s^{\prime}}$ commute. Conversely, given $\mathbf{H}_{t}$ and $\boldsymbol{\Sigma}_{t}$, the drift and diffusion coefficients can be obtained via:
@@
Therefore, theorem 1 allows one to obtain the linear diffusion process governing the evolution of $\mathbf{x_{t}}\mid\mathbf{x}_{0}$ defined through $\mathbf{H}_{t}$ and $\mathbf{\Sigma}_{t}$, which in turn makes training and sampling from continuous-time models possible.
@@
for $\alpha_{t},\beta_{t},\gamma_{t}$ being the null space drift, null space diffusion and range space diffusion coefficients respectively. We refer to the process resulting from eqs. 8a and 8b as the System-embedded Diffusion Bridge Model (SDB).
@@
$\mathbf{A}^{+}\mathbf{A}\mathbf{x}_{t}$ $=\mathbf{A}^{+}(\mathbf{A}\mathbf{x}_{0}+\gamma_{t}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon})=\mathbf{A}^{+}\mathbf{A}\mathbf{x}_{0}+\gamma_{t}\mathbf{A}^{+}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon},$ (9a)
$(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{t}$ $=\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{0}+\beta_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\boldsymbol{\epsilon}^{\prime},$ (9b)
@@
Applying theorem 1 to the coefficients from eq. 8 leads to the following drift and diffusion coefficients for SDB:
@@
[…]
@@
Additional technical details, including the chosen network parameterization and an extended discussion of nonlinear measurement systems, are provided in the Appendix.
@@
$\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ $\beta_{t}\mathbf{I}$ $\sigma^{2}(t)\mathbf{I}$ $g^{2}(t)\mathbf{I}$ $g^{2}(t)\mathbf{I}$ $\left(\frac{\text{d}\gamma_{t}}{\text{d}t}\right)\mathbf{A}^{+}\boldsymbol{\Sigma}\mathbf{A}^{+^{\top}}+$ $\left(\frac{\text{d}\beta_{t}}{\text{d}t}-2\beta_{t}\frac{\text{d}}{\text{d}t}\log{\alpha_{t}}\right)(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})$
$\boldsymbol{\mu}_{t}$ $\frac{\bar{\sigma}^{2}_{t}}{\bar{\sigma}^{2}_{t}+\sigma_{t}^{2}}\mathbf{x}_{0}+$ $\frac{\sigma^{2}_{t}}{\sigma^{2}_{t}+\sigma_{t}^{2}}\mathbf{x}_{1}$ $\mathbf{x}_{1}+$ $(\mathbf{x}_{0}-\mathbf{x}_{1})e^{-\bar{\theta}_{0:t}}$ $e^{-\bar{\theta}_{0:t}}\frac{\bar{\sigma}_{t:1}^{2}}{\bar{\sigma}_{0:1}^{2}}\mathbf{x}_{0}+$ $[(1-e^{-\bar{\theta}_{0:t}})\frac{\bar{\sigma}_{t:1}^{2}}{\bar{\sigma}_{0:1}^{2}}+$ $e^{-2\bar{\theta}_{t:1}}\frac{\bar{\sigma}_{0:t}^{2}}{\bar{\sigma}_{0:1}^{2}}]\mathbf{x}_{1}$ $\frac{\text{SNR}_{1}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{1}}\mathbf{x}_{1}+$ $\alpha_{t}(1-\frac{\text{SNR}_{1}}{\text{SNR}_{t}})\mathbf{x}_{0}$ $\mathbf{A}^{+}\mathbf{A}\mathbf{x}_{0}+\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{0}$
@@
5 Experiments
@@
We compare SDB with both supervised bridge methods and unsupervised plug-and-play diffusion-based baselines. For the former, we include I2SB (Liu et al., 2023), IR-SDE (Luo et al., 2023a), GOUB (Yue et al., 2024), and DDBM (Zhou et al., 2024). For the latter, we pick DPS (Chung et al., 2023a), $\Pi$GDM (Song et al., 2023a), and DDNM (Wang et al., 2023), all of which rely on the assumption of a noisy linear measurement model. Detailed descriptions of all baselines are provided in the Appendix. Unlike these methods, which rely on scalar SDEs, SDB introduces a more general matrix-valued formulation that allows structured control over range and null space components. We summarize the resulting conceptual differences to baseline bridge methods in table 1. Following standard evaluation practice (Luo et al., 2023a; Yue et al., 2024), we report perceptual scores (FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018)) and reconstruction metrics (PSNR, SSIM).
@@
We evaluate SDB on four inverse problems with varying measurement system complexities, using original images at a resolution of $256\times 256$. Building on prior work (Luo et al., 2023a; Yue et al., 2024), we first consider inpainting on CelebA-HQ (Karras et al., 2018), where $\mathbf{A}$ is a masking operator and $\boldsymbol{\Sigma}=\mathbf{0}_{m\times m}$ (noiseless). The measurements consist of the masked original images ($m=d$), which are equivalent to the PRs ($\mathbf{A}^{+}=\mathbf{I}$). Additionally, we follow previous studies by examining superresolution on DIV2K (Agustsson and Timofte, 2017; Timofte et al., 2017), where $\mathbf{A}$ represents a 4×4$\times$4 × downsampling operator ($d=4m$), implemented through average pooling. For this task, $\mathbf{A}^{+}$ reconstructs the images using nearest-neighbor interpolation. To evaluate more complex and practical inverse problems, we additionally propose two medical imaging tasks: CT reconstruction on the RSNA Intracranial Hemorrhage dataset (Anouk Stein et al., 2019) and MRI reconstruction on the Br35H dataset (Merlin, 2022), both using 2D axial brain scan slices.
@@
For MRI reconstruction, $\mathbf{y}$ consists of undersampled Fourier-domain measurements. The system matrix is modeled as $\mathbf{A}=\mathbf{M}\boldsymbol{\mathcal{F}}$, where $\boldsymbol{\mathcal{F}}$ is the Fourier transform and $\mathbf{M}$ is a masking matrix. We sample $\mathbf{M}$ such that $\lambda_{1}$ controls the percentage of the lowest frequencies to be kept (a deterministic operation), while $\lambda_{2}$ specifies the percentage of frequencies to sample from the remaining ones (a stochastic operation). The noisy setting is similarly modeled with a scalar covariance matrix $\boldsymbol{\Sigma}=\sigma_{2}^{2}\mathbf{I}$, where $\sigma_{2}^{2}=5.0$. We pick $\lambda_{1}=16$ and $\lambda_{2}=30$ as defaults.
5.1 General evaluation

### v2

@@ 3 Background
where $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}$ is the score function and $\overline{\mathbf{w}}_{t}$ is the Wiener process with reversed time. The forward-reverse relationship ensures that the marginal distributions $p(\mathbf{x}_{t})$, defined by eq.˜1 and eq.˜2, remain consistent.
In standard SGM for image synthesis, $p(\mathbf{x}_{1})$ is typically chosen as an isotropic Gaussian. To generate samples from $p(\mathbf{x}_{0})$, one samples from $p(\mathbf{x}_{1})$ and solves eq.˜2 using a time discretization scheme. This requires the unknown score function $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}$, which is generally approximated by a neural network $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ trained with the score-matching loss (Song et al., 2021b):
@@
The linear formulation in eq.˜1 and eq.˜2 is a special case of the broader SGM framework introduced by Song et al. (2021b), where drift and diffusion coefficients can be arbitrarily complex functions of $\mathbf{x}_{t}$ and $t$. This linear formulation, however, generalizes the well-established scalar case. Specifically, the Variance-Preserving (VP) SDE, which maintains constant variance for $\mathbf{x}_{t}$ across time, can be recovered with $\mathbf{F_{t}}=-\frac{1}{2}\beta_{t}\mathbf{I}$ and $\mathbf{G}_{t}=\sqrt{\beta_{t}}\mathbf{I}$. The Variance-Exploding (VE) SDE, which induces unbounded growth in the variance of $p(\mathbf{x}_{1})$ as $t\to 1$, is obtained with $\mathbf{F_{t}}=\mathbf{0}$ and $\mathbf{G}_{t}=\sqrt{\frac{\text{d}\sigma^{2}_{t}}{\text{d}t}}\mathbf{I}$. While these scalar matrices offer simplicity, they also raise the question of whether more sophisticated designs could yield additional benefits.
@@
where $\mathbf{y}\in\mathbb{R}^{m}$ is the measurement, $\boldsymbol{A}\in\mathbb{R}^{m\times d}$ is the system response matrix, $\boldsymbol{\epsilon}\sim\boldsymbol{\mathcal{N}}(\mathbf{0}_{m},\mathbf{I}_{m\times m})$ is the measurement noise with covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{m\times m}$ and $p(\mathbf{y}\mid\mathbf{x})=\boldsymbol{\mathcal{N}}(\mathbf{A}\mathbf{x},\boldsymbol{\Sigma})$ is the forward model. We refer to finding the unknown $\mathbf{x}$ based on the provided $\mathbf{y}$ as solving a linear inverse problem with additive Gaussian noise (Tarantola, 2005). This formulation is central to many practical applications, such as signal processing, medical imaging, and remote sensing (Mueller and Siltanen, 2012), particularly in the case of noninvertible linear systems, where either $m\neq d$ or $\mathbf{A}$ is rank-deficient.
@@
Consider the simplest noiseless ($\boldsymbol{\Sigma}=\mathbf{0}_{m\times m}$) setting of eq.˜4 with invertible $\mathbf{A}$. In this case, the original signal $\mathbf{x}$ can be fully recovered by simply using the inverse of $\mathbf{A}$, i.e., $\mathbf{x}=\mathbf{A}^{-1}\mathbf{y}$. However, this is only possible when $m=d$ and $\mathbf{A}$ is full-rank, which greatly limits the scope of potential inverse problems. In practice, the matrix $\mathbf{A}$ is typically not invertible, and noise arises naturally due to, e.g., imperfections in the measurement process.
@@
Many unsupervised approaches, assuming access to a specific noisy linear system (eq.˜4), offer tailored solutions that improve performance under additional assumptions (Song et al., 2023a; Wang et al., 2023; Chung et al., 2023a; Garber and Tirer, 2024). In contrast, state-of-the-art (SOTA) diffusion bridges (Liu et al., 2023; Luo et al., 2023a; Yue et al., 2024; Zhou et al., 2024) make no assumptions about the underlying system during training, focusing on general mappings between arbitrary distributions. This can be suboptimal when the system is (even approximately) known. For instance, in image inpainting, these methods do not distinguish between the known range (unmasked) and the missing null part (masked), resulting in redundant noise in the range space, which is expected to be noiseless. In more complex scenarios, such redundancy could reduce efficiency, while ignoring system-specific information may hinder generalization. Thus, developing a supervised diffusion bridge tailored to a specific system remains an important yet unresolved challenge.
@@ 4 Method
Following eq.˜4, it is clear that $\mathbf{y}\mid\mathbf{x}\sim\boldsymbol{\mathcal{N}}(\mathbf{A}\mathbf{x},\boldsymbol{\Sigma})$, i.e., for a given signal $\mathbf{x}$, a noisy linear system applies a linear transformation $\mathbf{A}$ and adds (possibly correlated) Gaussian noise to sample a measurement. On the other hand, given a clean sample $\mathbf{x}_{0}$, eq.˜1 transforms it to $\mathbf{x}_{t}$ by applying a cascade of such noisy linear systems, which is a noisy linear system itself. That is, $\mathbf{x}_{t}\mid\mathbf{x}_{0}\sim\boldsymbol{\mathcal{N}}(\mathbf{H}_{t}\mathbf{x}_{0},\boldsymbol{\Sigma}_{t})$ for some time-dependent matrices $\mathbf{H}_{t},\boldsymbol{\Sigma}_{t}$. Hence, one may equivalently arrive at $\mathbf{y}\mid\mathbf{x}$ (or some linear transformation of $\mathbf{y}$) through an application of a series of noisy linear systems to $\mathbf{x}$ instead of just a single one. In the limit, this would imply the existence of a general linear SDE that performs such mapping.
@@
(Tivnan et al., 2025) Assume that $\mathbf{x}_{t}\mid\mathbf{x}_{0}$ evolves according to the linear forward process from eq.˜1. Then, $\mathbf{x}_{t}\mid\mathbf{x}_{0}\sim\mathbf{\mathcal{N}}(\mathbf{H}_{t}\mathbf{x}_{0},\boldsymbol{\Sigma}_{t})$, where
@@
where $\boldsymbol{\Omega}_{t}$ is the Magnus expansion and the approximation becomes equality if, for all $s,s^{\prime}\in[0,t]$, $\mathbf{F}_{s},\mathbf{F}_{s^{\prime}}]=0$, i.e., $\mathbf{F}_{s}$ and $\mathbf{F}_{s^{\prime}}$ commute. Conversely, given $\mathbf{H}_{t}$ and $\boldsymbol{\Sigma}_{t}$, the drift and diffusion coefficients can be obtained via:
@@
Therefore, theorem˜1 allows one to obtain the linear diffusion process governing the evolution of $\mathbf{x_{t}}\mid\mathbf{x}_{0}$ defined through $\mathbf{H}_{t}$ and $\mathbf{\Sigma}_{t}$, which in turn makes training and sampling from continuous-time models possible.
@@
for $\alpha_{t},\beta_{t},\gamma_{t}$ being the null space drift, null space diffusion and range space diffusion coefficients respectively. We refer to the process resulting from eqs.˜8a and 8b as the System-embedded Diffusion Bridge Model (SDB).
@@
$\mathbf{A}^{+}\mathbf{A}\mathbf{x}_{t}$ $=\mathbf{A}^{+}(\mathbf{A}\mathbf{x}_{0}+\gamma_{t}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon})=\mathbf{A}^{+}\mathbf{A}\mathbf{x}_{0}+\sqrt{\gamma_{t}}\mathbf{A}^{+}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon},$ (9a)
$(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{t}$ $=\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{0}+\sqrt{\beta_{t}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\boldsymbol{\epsilon}^{\prime},$ (9b)
@@
Applying theorem˜1 to the coefficients from eq.˜8 leads to the following drift and diffusion coefficients for SDB:
@@
For the linear SDE defined in eq.˜1 with $\mathbf{F}_{t}$ and $\mathbf{G}_{t}$ given by eqs.˜10a and 10b respectively, let $\alpha_{t}=\frac{\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}},\beta_{t}=\frac{\sigma_{t}^{2}\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}$ for $\sigma_{t}^{2}=\int_{0}^{t}g^{2}(\tau)d\tau,\bar{\sigma}_{t}^{2}=\int_{t}^{1}g^{2}(\tau)d\tau$ for some non-negative function $g(t)$. Assuming that $g(t)\rightarrow 0$ uniformly for all $t$, the null space part of this SDE reduces to an OT-ODE:
@@
Principled posterior sampling.
In the asymptotic limit of infinite data and model capacity, it is natural to ask whether SDB constitutes an exact probabilistic model of the underlying inverse problem—that is, whether it produces principled posterior samples from $p(\mathbf{x}|\mathbf{y})$. The following result provides a positive answer under mild conditions on the forward process coefficients, with the proof deferred to the Appendix.
Theorem 3
Let the forward measurement model be $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(\mathbf{A}\mathbf{x},\boldsymbol{\Sigma})$. Then, under the SDB dynamics defined by eq.˜1, eq.˜7, and eq.˜8, with time-dependent scalar coefficients $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$ satisfying
$\lim_{t\rightarrow 1}\gamma_{t}=1,\quad\lim_{t\rightarrow 1}\frac{\alpha_{t}^{2}}{\beta_{t}}=0,$
the corresponding reverse-time SDE generates asymptotically exact samples from the Bayesian posterior distribution $p(\mathbf{x}|\mathbf{y})$.
@@
$\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ $\beta_{t}\mathbf{I}$ $\sigma^{2}(t)\mathbf{I}$ $g^{2}(t)\mathbf{I}$ $g^{2}(t)\mathbf{I}$ $\frac{\text{d}\gamma_{t}}{dt}\mathbf{A}^{+}\boldsymbol{\Sigma}\mathbf{A}^{+^{\top}}+$ $\left(\frac{\text{d}\beta_{t}}{\text{d}t}-2\beta_{t}\frac{\text{d}}{\text{d}t}\log{\alpha_{t}}\right)(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})$
$\boldsymbol{\mu}_{t}$ $\frac{\bar{\sigma}^{2}_{t}}{\bar{\sigma}^{2}_{t}+\sigma_{t}^{2}}\mathbf{x}_{0}+$ $\frac{\sigma^{2}_{t}}{\sigma^{2}_{t}+\sigma_{t}^{2}}\mathbf{x}_{1}$ $\mathbf{x}_{1}+$ $(\mathbf{x}_{0}-\mathbf{x}_{1})e^{-\bar{\theta}_{0:t}}$ $e^{-\bar{\theta}_{0:t}}\frac{\bar{\sigma}_{t:1}^{2}}{\bar{\sigma}_{0:1}^{2}}\mathbf{x}_{0}+$ $(1-e^{-\bar{\theta}_{0:t}})\frac{\bar{\sigma}_{t:1}^{2}}{\bar{\sigma}_{0:1}^{2}}+$ $e^{-2\bar{\theta}_{t:1}}\frac{\bar{\sigma}_{0:t}^{2}}{\bar{\sigma}_{0:1}^{2}}]\mathbf{x}_{1}$ $\frac{\text{SNR}_{1}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{1}}\mathbf{x}_{1}+$ $\alpha_{t}(1-\frac{\text{SNR}_{1}}{\text{SNR}_{t}})\mathbf{x}_{0}$ $\mathbf{A}^{+}\mathbf{A}\mathbf{x}_{0}+\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{0}$
@@
Hyperparameters $\beta_{t},$ $\sigma_{t}^{2}=\int_{0}^{t}{g^{2}(\tau)d\tau},$ $\bar{\sigma}_{t}^{2}=\int_{t}^{1}{g^{2}(\tau)d\tau}$ $\theta_{t}=\frac{\sigma^{2}_{t}}{\lambda^{2}},$ $\bar{\theta}_{t}=\int_{0}^{t}{\theta_{\tau}}d\tau$ $\theta_{t}=\frac{g^{2}(t)}{2\lambda^{2}},$ $\bar{\theta}_{s:t}=\int_{s}^{t}\theta_{\tau}d\tau,$ $\bar{\sigma}_{s:t}^{2}=\frac{g^{2}(t)}{2\theta_{t}}(1-e^{-2\bar{\theta}_{s:t}})$ $\text{SNR}_{t}=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}},$ $\alpha_{t}=\exp{\left(-\frac{1}{2}\int_{0}^{t}{g^{2}(\tau)d\tau}\right)},$ $\sigma_{t}^{2}=1-\alpha_{t}^{2}$ $\alpha_{t},\beta_{t},\gamma_{t}$ follow from theorem˜2
5 Experiments
Algorithm 1 SDB Training
@@ 5 Experiments
2:for each iteration do
3:  $\mathbf{x}\sim p(\mathbf{x}_{0})$ $\triangleright$ sample clean data $\mathbf{x}$
4:  $t\sim p(t)$ $\triangleright$ sample diffusion timestep $t$
5:  $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{m}$ $\triangleright$ sample range-space Gaussian noise $\boldsymbol{\epsilon}$
6:  $\boldsymbol{\epsilon}^{\prime}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{d}$ $\triangleright$ sample null-space Gaussian noise $\boldsymbol{\epsilon}^{\prime}$
7:  $\mathbf{x}_{t}\leftarrow[\mathbf{A}^{+}\mathbf{A}+\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})]\mathbf{x}+\gamma_{t}^{\frac{1}{2}}\mathbf{A}^{+}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon}+\beta_{t}^{\frac{1}{2}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\boldsymbol{\epsilon}^{\prime}$ $\triangleright$ forward step
8:  $L_{\theta}\leftarrow\left\|\mathcal{\boldsymbol{D}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}\right\|_{1}$ $\triangleright$ compute reconstruction loss for $\mathbf{x}$
9:  $\theta\leftarrow\text{optimizer}(\nabla_{\theta}L_{\theta})$ $\triangleright$ update network parameters $\theta$
10:end for
11:return $\theta$
Algorithm 2 SDB Sampling (Euler-Maruyama)
[…]
7:for $i\in\{1,\ldots,N\}$ do
8:  $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{m}$ $\triangleright$ sample range-space noise
9:  $\boldsymbol{\epsilon}^{\prime}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{d}$ $\triangleright$ sample null-space noise
10:   $\mathbf{x}_{t-\Delta t}\leftarrow\mathbf{x}_{t}+\Delta t\left[(f_{t}\mathbf{I}-2\mathbf{F}_{t})(\mathbf{H}_{t}\mathcal{\boldsymbol{D}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}_{t})-\mathbf{F}_{t}\mathbf{x}_{t}\right]+$ $\Delta t^{1/2}\left[\left(\frac{\text{d}\gamma_{t}}{\text{d}t}\right)^{\frac{1}{2}}\mathbf{A}^{+}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon}+\left(\frac{\text{d}\beta_{t}}{\text{d}t}-2\beta_{t}\frac{\text{d}}{\text{d}t}\log{\alpha_{t}}\right)^{\frac{1}{2}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\boldsymbol{\epsilon}^{\prime}\right]$ $\triangleright$ update
11:  $t\leftarrow t-\Delta t$ $\triangleright$ decrement time
12:end for
13:return $\mathbf{x}_{0}$ $\triangleright$ final sample
@@
[…]
@@
We evaluate SDB on four inverse problems with varying measurement system complexities, using original images at a resolution of $256\times 256$. Building on prior work (Luo et al., 2023a; Yue et al., 2024), we first consider inpainting on CelebA-HQ (Karras et al., 2018), where $\mathbf{A}$ is a masking operator and $\boldsymbol{\Sigma}=\mathbf{0}_{m\times m}$ (noiseless). The measurements consist of the masked original images ($m=d$), which are equivalent to the PRs ($\mathbf{A}^{+}=\mathbf{I}$). Additionally, we follow previous studies by examining superresolution on DIV2K (Agustsson and Timofte, 2017; Timofte et al., 2017), where $\mathbf{A}$ represents a $4\times$ downsampling operator ($d=4m$), implemented through average pooling. For this task, $\mathbf{A}^{+}$ reconstructs the images using nearest-neighbor interpolation. To evaluate more complex and practical inverse problems, we additionally propose two medical imaging tasks: CT reconstruction on the RSNA Intracranial Hemorrhage dataset (Anouk Stein et al., 2019) and MRI reconstruction on the Br35H dataset (Merlin, 2022), both using 2D axial brain scan slices.
@@
[…]